# Federated Automated Feature Engineering

### v1

@@ 1 Abstract
Automated feature engineering (AutoFE) is used to automatically create new features from original features to improve predictive performance without needing significant human intervention and domain expertise. Many algorithms exist for AutoFE, but very few approaches exist for the federated learning (FL) setting where data is gathered across many clients and is not shared between clients or a central server. We introduce AutoFE algorithms for the horizontal, vertical, and hybrid FL settings, which differ in how the data is gathered across clients. To the best of our knowledge, we are the first to develop AutoFE algorithms for the horizontal and hybrid FL cases, and we show that the downstream model performance of federated AutoFE is similar to the case where data is held centrally and AutoFE is performed centrally.
@@ 2 Introduction
This paper focuses on automated feature engineering in the context of federated learning, where clients are not able to send data to the server, but we still build informative new features from the original raw features. We introduce algorithms for all three FL settings: horizontal, vertical, and hybrid. The horizontal algorithm allows each client to perform any AutoFE method of choice, the server aggregates the engineered feature strings, and then uses a communication-minimizing, hyperband-inspired selection method [hyperband] to determine which engineered features to keep. The hybrid and vertical FL methods hinge on homomorphic encryption to safely form combinations of features that reside on different clients, then use a similar selection scheme as the horizontal algorithm to determine the best engineered features to keep. All of the algorithms are iterative in that they gradually construct more and more complicated engineered features each iteration.
@@
    To the best of our knowledge, we are the first to develop federated AutoFE algorithms in the horizontal and hybrid FL settings. We also provide a new vertical AutoFE algorithm.
@@
    We provide experiments demonstrating that our horizontal FL AutoFE algorithm performs similarly to the case where AutoFE is performed centrally. This is a major achievement in FL where typically federated algorithms perform notably worse than the centralized counterparts.
In Section 3, we review the relevant work on both federated learning and AutoFE and discuss the minimal work that exist on AutoFE in the vertical FL setting. In Section 4, we detail our proposed AutoFE algorithms for the horizontal, vertical, and hybrid FL settings. In Section 5, we show results of our algorithm in the horizontal FL setting and demonstrate the performance is comparable to the case where data is located centrally and AutoFE is performed centrally.
@@ 3 Related Work
Despite the great deal of work in federated learning and automated feature engineering, there is relatively limited work on automated feature engineering in the federated learning setting. All of the existing work at the time of this writing focused on the vertical case, where each client has different features, and all clients have the same samples. There exists work in AutoML in a federated setting, however automated feature engineering is not one of the aspects discussed in this work [autofl]. FLFE is an algorithm for vertical federated AutoFE that uses quantile sketching to encode features and uses deep neural networks to predict if the transformation of features is useful [vertfefl]. The neural networks are trained on a large quantity of datasets. Finally, homomorphic encryption is used when constructing the best candidate features for building the final model. However, this work shows only small gains compared to using only the raw features, does not compare to AutoFE methods in the centralized setting, and does not provide any source code making reproducibility very difficult. Another recent work uses chi-square binning and homomorphic encryption to tackle the vertical federated AutoFE setting [vertfefl2]. However, this work does not share source code, only tests on two datasets, and only shares runtimes and not the actual test scores of using the new features. We believe this makes the work very difficult to reproduce and compare. It is clear that there is a large research gap in all three settings of federated AutoFE: horizontal, vertical, and hybrid. Our work provides algorithms for all of these settings.
@@ 4 Algorithms
The most straightforward approach to performing AutoFE in the federated learning setting is to simply run the AutoFE algorithm as usual but perform all model trainings and evaluations using federated algorithms. However, since most AutoFE algorithms require thousands of model trainings and evaluations and each of those model trainings would require many communications between clients and the server, the communication overhead makes this naive approach infeasible. The development of our federated AutoFE algorithms for each FL setting rely on minimizing this communication overhead.

### v2

@@ 1 Abstract
Automated feature engineering (AutoFE) is used to automatically create new features from original features to improve predictive performance without needing significant human intervention and domain expertise. Many algorithms exist for AutoFE, but very few approaches exist for the federated learning (FL) setting where data is gathered across many clients and is not shared between clients or a central server. We introduce AutoFE algorithms for the horizontal, vertical, and hybrid FL settings, which differ in how the data is gathered across clients. To the best of our knowledge, we are the first to develop AutoFE algorithms for the horizontal and hybrid FL cases, and we show that the downstream test scores of our federated AutoFE algorithms is close in performance to the case where data is held centrally and AutoFE is performed centrally.
@@ 2 Introduction
This paper focuses on automated feature engineering in the context of federated learning, where clients are not able to send data to the server, but we still build informative new features from the original raw features. We introduce algorithms for all three FL settings: horizontal, vertical, and hybrid. The horizontal algorithm we propose extends IIFE [iife] to the federated setting, which we term Fed-IIFE. The hybrid and vertical FL methods hinge on homomorphic encryption and differential privacy to safely form combinations of features that reside on different clients, along with a feature selection scheme to determine which features to keep. All of the algorithms are iterative in that they gradually construct more and more complicated engineered features each iteration.
@@
    To the best of our knowledge, we are the first to develop federated AutoFE algorithms in the horizontal and hybrid FL settings. We also provide a new vertical AutoFE algorithm that is safer from client privacy leaks than the one existing approach in this setting.
@@
    We provide experiments demonstrating that our horizontal FL AutoFE algorithm performs similarly to the case where AutoFE is performed centrally, even in cases where the data distributions across clients are highly non-IID. We validate the ability of federated interaction information to determine synergy between feature pairs with a synthetic experiment. We also demonstrate that the important engineered features generated by centralized IIFE have similarly generated features in Fed-IIFE, even in highly non-IID settings.
  3. 3.
    We demonstrate that our hybrid and vertical AutoFE algorithms still show significant improvements over the baseline (no engineered features) even with the strictly limiting requirements of homomorphic encryption and differential privacy.
In Section 3, we review the relevant work on both federated learning and AutoFE and discuss the minimal and incomplete work that exists on AutoFE in the vertical FL setting. In Section 4, we detail our proposed AutoFE algorithms for the horizontal, vertical, and hybrid FL settings. In Section 5, we show experimental results for all three FL settings, and verify key aspects of Fed-IIFE such as the validity of federated interaction information and similarity of engineered features between the centralized and FL settings.
@@ 3 Related Work
Despite the great deal of work in federated learning and automated feature engineering, there is relatively limited work on automated feature engineering in the federated learning setting. All of the existing work at the time of this writing focuses on the vertical case, where each client has different features, and all clients have the same samples. There exists work in AutoML in a federated setting, however automated feature engineering is not one of the aspects discussed in this work [autofl]. FLFE is an algorithm for vertical federated AutoFE that uses quantile sketching to encode features and uses deep neural networks to predict if a particular transformation of features is useful [vertfefl]. The neural networks are trained on a number of datasets, but these datasets are not shared, so it is impossible to reproduce their results. Since these datasets are actually used to build the feature engineering framework and not just used for evaluating the algorithm, it is necessary to have these datasets to properly evaluate the algorithm. Finally, homomorphic encryption is used when constructing the best candidate features for building the final model. However, this work shows only small gains compared to using only the raw features, does not compare to AutoFE methods in the centralized setting, and does not provide any source code making reproducibility very difficult. Furthermore, clients may be able to reconstruct the features of other clients from the decrypted engineered features; we address this in our work by utilizing differential privacy. We believe this makes the work very difficult to reproduce and compare. It is clear that there is a large research gap in all three settings of federated AutoFE: horizontal, vertical, and hybrid. Our work provides algorithms for all of these settings.
@@ 4 Algorithms
Adapting existing AutoFE algorithms to the horizontal FL setting is not straightforward. Many of these algorithms rely on expanding a very large number of engineered features and evaluating each of these new features, such as OpenFE, EAAFE, and AutoFeat. These extreme number of FL evaluations result in too much communication overhead to be practical. Furthermore, many algorithms have operations that are not straightforward to extend to the FL case, such as the LightGBM boosting evaluations that OpenFE performs.
An alternative to directly incorporating FL into the algorithm is to perform AutoFE on each client separately, send the feature transformation strings to the central server where it takes the union of these transformations, and then send this union back to the clients. An optional step of performing federated feature selection after finding this union can further help performance; we provide such a federated feature selection algorithm in Section 7.3 of the Appendix. However, we find that this approach does not perform well because the feature engineering step does not incorporate any information from the other clients. This performance degradation is especially apparent when data is distributed across clients in a non-IID manner. We discuss this further in Section 7.1.2 of the Appendix.
Fortunately, extending IIFE to the FL setting is relatively straightforward and results in good empirical performance compared to centralized IIFE, which is one of the top-performing AutoFE algorithms at the time of this writing [iife]. We detail this algorithm, Fed-IIFE, in Section 4.1. The cases of vertical and hybrid FL are more complicated than horizontal FL and are discussed in Section 4.2.